# Simulating stochastic population dynamics: The Linear Noise Approximation can capture non-linear phenomena

### v3

@@ Abstract
There is an abundance of highly stochastic, non-linear phenomena present in population dynamics across various fields, including molecular biology, epidemiology, and ecology. Oscillations and multi-stability are particularly prevalent in gene regulation systems. None of the currently available stochastic models for population dynamics are accurate and computationally efficient for long-term predictions. A prominent model in this field, known as the Linear Noise Approximation (LNA), is computationally efficient for tasks such as simulation, sensitivity analysis, and parameter estimation, still, it is only accurate for linear systems and short-time predictions. Other models may achieve greater accuracy across a broader range of systems, but they sacrifice computational efficiency and analytical tractability. This paper demonstrates that, with specific modifications, the LNA can indeed accurately capture non-linear dynamics in population dynamics. We introduce a new framework that employs centre manifold theory, a classical concept in non-linear dynamical systems. This approach allows us to identify simple, specialised modifications to the LNA tailored to classes of qualitatively similar non-linear dynamical systems. With these modifications, the LNA can achieve accurate long-term predictions without compromising its computational efficiency. We apply our methodology to a class of oscillatory systems and a class of bi-stable systems, and we provide multiple examples in molecular population dynamics that showcase accurate long-term predictions, and significant improvements in computational efficiency.
@@ 1 Introduction
In molecular biology, ecology, epidemiology and other fields, there is a plethora of non-linear dynamical phenomena such as oscillations and multi-stabilities. Examples include the circadian rhythms [21], embryonic development oscillations [39], cell signalling oscillations [4], multi-stabilities in cell cycle [45, 46, 2], cell differentiation [8, 19], and apoptosis [3], the central carbon metabolism in E.coli [26], predator-prey oscillations in ecology [11], and epidemic oscillations [62]. These non-linear dynamics are caused by feed-forward loops [38], negative feedback loops [38, 48] and other mechanisms, which are operated by networks of interacting species. For example, gene expression oscillations are often generated when a gene directly suppresses its own expression or activates the expression of its suppressors. In infectious diseases outbreaks, compartments of the population, such as infected and susceptible individuals, interact between each other and phenomena like awareness and fatigue can generate oscillatory behaviors [62]. The effects of stochasticity in these interactions between different populations, caused by molecular diffusion or, more generally, the complex environment in which these interactions occur, are significant, especially when dealing with smaller populations. Yet modeling the stochasticity using fast, accurate, scalable methods is an unsolved problem.
@@
These networks of stochastic interactions between different populations are often called reaction networks. The term reaction is not only used for biochemical reactions but any event that directly causes a change in the population of one or more species or groups. The stochastic dynamics of reaction networks, can be described, under certain assumptions, by Markov processes $\{{\bm{Y}}(t)=(Y_{1}(t),\dots,Y_{n}(t))^{T}:t\geq 0\}$, describing the evolution of $n$ different populations over time. The so-called (chemical) master equation [13], which describes the evolution of the probability distribution of the species populations, can be solved only rarely, severely limiting its tractability. The stochastic simulation algorithm (SSA) [14] produces exact simulations of the master equation but it is far too slow to be used in most applied settings since it simulates every single reaction event.
@@
At present, there are several methods available for approximating the master equation, which aim to accelerate simulations. Notable examples include tau leaping algorithms [17, 18] and the numerical integration of the (chemical) Langevin equations [16]. Both of these approaches significantly increase simulation speed; however, they leave no scope for the analytical study of the probability distributions of the populations and can be extremely slow for sensitivity analysis and parameter estimation. Other recent methods [6, 28] focus on simplifying protein-binding reactions, which frequently appear in molecular biology networks.
@@
Much work has followed from van Kampen’s system size expansion [58], which uses a system size parameter $\Omega$ (like cell volume in molecular biology, maximum population size in ecology or epidemiology) to derive approximations of the master equation. Under certain conditions, the deterministic process $\{{\bm{x}}(t):t\geq 0\}$ which solves the macroscopic reaction-rate equations was shown to be the limit, as $\Omega\to\infty$, of the solutions of the master equation [32, 33]. The Linear Noise Approximation (LNA), derived using the system size expansion, is a stochastic model that provides even more rapid simulation than the aforementioned stochastic methods as well as being the only model in this field that allows analytical expressions of probability distributions of populations at arbitrary time-points. This makes the LNA a powerful tool for sensitivity analysis and statistical inference [41, 42, 54, 43, 31, 10, 20, 49].
@@
However, the important drawback of using the LNA is its inability to approximate the long-term behavior of non-linear population dynamics. Indeed, when the system dynamics are non-linear [61], predictions of the state of the system at time $t$ given the state at an earlier time $s$ become inaccurate as tst-s becomes large. For a given system, we can qualitatively examine the dynamics of its paths and identify different phases in the populations’ evolution - for instance their peaks and troughs. If we were to examine the distribution of entire stochastic paths generated by the SSA in a finite number of time-points then we find that, the phases of such paths become out of sync with what the LNA predicts their phases to be. Hence the LNA makes incorrect predictions of the dynamics thereafter. In [41], it was shown that in the specific setting where the macroscopic ($\Omega\to\infty$) deterministic approximation has an attractive limit cycle solution, correcting these drifts in phase predicted by the LNA is sufficient to produce simulations that are nearly identical to those produced by the SSA. This approach enables the construction of both fast and accurate algorithms for stochastic simulation, sensitivity analysis, and statistical inference of this class of non-linear systems [41].
There are three main results in this paper. Firstly, we show that the analytically tractable LNA model, which is known to be long-time accurate only for linear systems, can also be applied to non-linear systems after appropriate modifications. In particular, we illustrate that it can accurately describe long-term stochastic dynamics for arguably the two most abundant non-linear phenomena: oscillations and bi-stability. We demonstrate this on several systems including the oscillatory systems in [35, 63, 4], and bi-stable systems in [12, 2, 19]. Secondly, as part of this effort, we develop a novel framework for constructing LNA-based models for non-linear dynamical systems. We adopt the term phase-corrected LNA (pcLNA)—originally introduced in [41]—to describe this stochastic modelling approach, which evolves the system using LNA equations while periodically adjusting the phase to mitigate phase drift. Fundamental to the pcLNA approach is a mapping G{G} that computes the phase of any stochastic state allowing for any phase drift to be corrected. We show that for the very large class of systems whose macroscopic reaction rate equation exhibits a non-hyperbolic equilibrium, the theory of centre manifolds of dynamical systems can be employed to explicitly define the phase correction map G{G}. Thirdly, we illustrate how the pcLNA framework can be applied to accurately describe reaction networks whose macroscopic deterministic evolution $\{{\bm{x}}(t):t\geq 0\}$ exhibits Hopf bifurcation and those exhibiting bi-stability in a region of their parameter space (see Figure 1). That is, systems with oscillations; sustained or non-sustained and often determined by only a small change in a parameter, or systems which can switch between two steady states due to stochasticity. We demonstrate that the pcLNA model for these systems offers accurate approximations of the master equation for long-term simulation whilst substantially reducing the simulation time by several orders of magnitude. A key implication of this accuracy is that it enables the tractability of pcLNA models to be harnessed for computationally intensive tasks such as sensitivity analysis, inference, and prediction.
@@
While the approach developed in [41] demonstrated that phase correction significantly improves the long-term accuracy of the LNA in oscillatory systems with stable limit cycles, it was limited to this specific dynamics. In contrast, the framework introduced in this paper generalises the concept of phase correction by introducing centre manifold theory to define and correct phase drifts to a broad class of systems only assumed to have a non-hyperbolic equilibrium, which includes the abundant classes of oscillatory and bi-stable systems. This theoretical foundation enables us to construct a unified, algorithmic approach for defining and correcting phase across diverse nonlinear regimes, allowing the pcLNA to extend the analytical tractability and computational efficiency of the LNA to a class of systems found in a wide range of scientific subjects.
@@
We begin in Section 2 by introducing the pcLNA modelling framework. We first review reaction networks and the system size expansion in Sections 2.1 and 2.2, before introducing the Linear Noise Approximation (LNA) in Section 2.3. We then discuss the LNA’s loss of accuracy in non-linear systems and introduce the concept of phase in this context. In Section 2.4, we present the general phase corrected LNA (pcLNA), followed by a discussion on how to define the phase correction map-using centre manifold theory—in Section 2.5. In Section 3, we apply our framework to reaction networks whose macroscopic reaction rate equation undergoes a Hopf bifurcation and develop the pcLNA algorithm specific to these systems. Section 4 focuses on bi-stable systems and constructs the corresponding pcLNA algorithms. We validate our framework through numerical investigations in Section 5, and conclude with a discussion of our findings and future directions in Section 6. Additional numerical investigations demonstrating the accuracy of the pcLNA modelling techniques on other reaction networks, as well as supporting technical results and derivations, are provided in the Supporting Information (SI).
@@
2 The pcLNA modelling framework
2.1 Population dynamics
We begin by presenting a mathematical formulation of population dynamics evolving through reaction networks. We describe this framework within the context of molecular biology, although analogous terminology can be applied in other fields such as epidemiology and ecology (see, for instance, [20, 64] and Figure 2).
A system of multiple different molecular sub-populations, $M_{1},M_{2},\dots,M_{n}$ has state vector, ${\bm{Y}}(t)=(Y_{1}(t),\ldots,Y_{n}(t))^{T}$ where $Y_{i}(t)\geq 0$, $i\in\{1,\ldots,n\}$, denotes the number of $M_{i}$ molecules at time $t$. These molecules undergo reactions $R_{j}$, $j\in\{1,\dots,r\}$, such as transcription, translation, degradation and translocation, which change the number of molecules of each specie,
@@ 2.1 Population dynamics
Here, the non-negative coefficients $k_{ij},k_{ij}^{\prime}\geq 0$ denote the number of molecules of $M_{i}$ involved as reactants and products, respectively, in reaction $R_{j}$. If $R_{j}$ occurs at time $t\geq 0$, ${\bm{Y}}(t)$ jumps to a new state ${\bm{Y}}(t)+{\bm{\nu}}_{j}$, where the transition vectors ${\bm{\nu}}_{j}=(k_{1j}^{\prime}-k_{1j},\dots,k_{nj}^{\prime}-k_{nj})^{T}\in\mathbb{R}^{n}$ are often called stoichiometric. Under the assumption that these reactions occur in a well-mixed solution [14], we can derive the propensity functions $\pi_{j}:\mathbb{R}^{n}\to[0,\infty)$ for each $j\in\{1,\dots,r\}$ with the probability of one $R_{j}$ reaction occurring in the next infinitesimal time interval $[t,t+dt)$
@@
The propensity functions $\pi_{j}$ along with the stoichiometric vector ${\bm{\nu}}_{j}$, $j=1,\dots,r$ completely specify the evolution of the reaction network.
The general form of the propensity function of the reaction in (1) is $\pi_{j}({\bm{y}})=c_{j}\prod_{i}y_{i}^{k_{ij}},$ with uni-molecular ($M_{i}\stackrel{{\scriptstyle c_{j}}}{{\longrightarrow}}\;\cdot$) and bi-molecular reactions ($M_{i}+M_{i^{\prime}}\stackrel{{\scriptstyle c_{j}}}{{\longrightarrow}}\;\cdot$) with propensity of the form $c_{j}y_{i}$ and $c_{j}y_{i}y_{i^{\prime}}$, respectively, being the most common type [60]. Here $c_{j}\in\mathbb{R}$ is the reaction rate constant corresponding to reaction $R_{j}$. The propensity functions are not restricted to the above general form. For instance, for enzymatic reactions involving species $M_{i}$ the propensity has the Michaelis-Menten form [29], $c_{2}y_{i}/(c_{1}+y_{i})$, while the Hill form [24], $c_{2}y_{i}^{h}/(c_{1}^{h}+y_{i}^{h})$, is also often used to reflect cooperative binding. For a discussion on the latter forms of propensities see [6].
These conditions give rise to a time-inhomogeneous, discrete state-space Markov process $\{{\bm{Y}}(t):t\geq 0\}$. The Kolmogorov forward equation [15] describing the evolution of the probability distribution, $P({\bm{y}},t):=P({\bm{Y}}(t)={\bm{y}}\mid{\bm{Y}}(0)={\bm{y}}_{0})$, of the stochastic process $\{{\bm{Y}}(t):t\geq 0\}$ is,
$\frac{\partial P({\bm{y}},t)}{\partial t}=\sum_{j=1}^{r}\pi_{j}({\bm{y}}-{\bm{\nu}}_{j})P({\bm{y}}-{\bm{\nu}}_{j},t)-\sum_{j=1}^{r}\pi_{j}({\bm{y}})P({\bm{y}},t).$ (3)
The Kolmogorov equation is often referred as (chemical) master equation. Despite the equation only being analytically tractable for few cases, an exact simulation algorithm is available, namely Gillespie’s stochastic simulation algorithm (SSA) [14]. We provide details of this in SI (Section S1).
Example (SIR model):
The SIR model used to describe population dynamics in infectious disease outbreaks is an example of a reaction network. We formulate it here to provide a concrete example of our general formulation above (see also Figure 2(A-B) for a summary). In the SIR network, there are three populations, say $M_{1},M_{2},M_{3}$, corresponding to the Susceptible (S), Infected (I), and Recovered (R) individuals. Two reactions, that is infection and recovery, say $R_{1}$ and $R_{2}$, can take place and can be described by
@@
where here $c_{1}$ and $c_{2}$ are the reaction rate constants for the two reactions. The state vector ${\bm{Y}}=(Y_{1},Y_{2},Y_{3})$ corresponds to the three populations $M_{1}$ (S), $M_{2}$ (I), and $M_{3}$ (R). The propensity functions are
@@
and the corresponding stoichiometry vectors ${\bm{v}}_{1}=(-1,+1,0)$ and ${\bm{v}}_{2}=(0,-1,+1)$. This can be used to formulate Equation (3) for this system.

### v4

@@ Abstract
Population dynamics in fields such as molecular biology, epidemiology, and ecology exhibit highly stochastic and non-linear behavior. In gene regulatory systems in particular, oscillations and multistability are especially common. Despite this, none of the currently available stochastic models for population dynamics are both accurate and computationally efficient for long-term predictions. A prominent model in this field, the Linear Noise Approximation (LNA), is computationally efficient for tasks such as simulation, sensitivity analysis, and parameter estimation; however, it is only accurate for linear systems and short-time predictions. Other models may achieve greater accuracy across a broader range of systems, but they sacrifice computational efficiency and analytical tractability. This paper demonstrates that, with specific modifications, the LNA can accurately capture non-linear dynamics in population processes. We introduce a new framework based on centre manifold theory, a classical concept from non-linear dynamical systems. This approach enables the identification of simple, system-specific modifications to the LNA, tailored to classes of qualitatively similar non-linear dynamical systems. With these modifications, the LNA can achieve accurate long-term simulations without compromising computational efficiency. We apply our methodology to classes of oscillatory and bi-stable systems, and present multiple examples from molecular population dynamics that demonstrate accurate long-term simulations alongside significant improvements in computational efficiency.
@@ 1 Introduction
In molecular biology, ecology, epidemiology and other fields, there is a multitude of non-linear dynamical phenomena such as oscillations and multi-stabilities. Examples include circadian rhythms [23], embryonic development oscillations [43], cell signalling oscillations [4], multi-stabilities in the cell cycle [50, 52, 2], cell differentiation [9, 21], and apoptosis [3], the central carbon metabolism in E.coli [28], predator-prey oscillations in ecology [13], and epidemic oscillations [70]. These non-linear dynamics are caused by feed-forward loops [42], negative feedback loops [42, 54] and other mechanisms, which are operated by networks of interacting species. For example, gene expression oscillations are often generated when a gene directly suppresses its own expression or activates the expression of its suppressors. In infectious diseases outbreaks, compartments of the population, such as infected and susceptible individuals, interact between each other and phenomena like awareness and fatigue can generate oscillatory behaviors [70]. The effects of stochasticity in interactions between different populations—arising from molecular diffusion or, more generally, from the complex environments in which these interactions occur—are significant, particularly for small populations. Yet modeling this stochasticity using fast, accurate, and scalable methods remains an unsolved problem.
@@
The networks of stochastic interactions between different populations are often called reaction networks. The term reaction is not only used for biochemical reactions but any event that directly causes a change in the population of one or more species or groups. The stochastic dynamics of reaction networks, can be described, under certain assumptions, by a Markov process $\{{\bm{Y}}(t)=(Y_{1}(t),\dots,Y_{n}(t))^{\top}:t\geq 0\}$, describing the evolution of $n$ different populations over time. The so-called (chemical) master equation [16], which describes the evolution of the probability distribution of the species populations over time, is analytically solvable only in rare cases, severely limiting its tractability. The stochastic simulation algorithm (SSA), also known as the Gillespie algorithm[17], produces exact simulations of the Markov process $\{{\bm{Y}}(t):t\geq 0\}$ but it is far too slow to be used in most applied settings since it simulates every single reaction event.
@@
At present, there are several methods available for approximating the master equation, which aim to accelerate simulations. Notable examples include tau leaping algorithms [20, 15] and the numerical integration of the chemical Langevin equations [19]. Both of these approaches significantly increase simulation speed; however, they leave no scope for the analytical study of the probability distributions of the populations and can be extremely slow for sensitivity analysis and parameter estimation. Other recent methods [7, 30] focus on simplifying protein-binding reactions, which frequently appear in molecular biology networks.
@@
Much work has followed from van Kampen’s system size expansion [66], which uses a system size parameter $\Omega$ (like cell volume in molecular biology, maximum population size in ecology or epidemiology) to derive approximations of the master equation. Under certain conditions, the deterministic process $\{{\bm{x}}(t)=(x_{1}(t),\dots,x_{n}(t))^{\top}:t\geq 0\}$ which solves the macroscopic reaction-rate equations was shown to be the law of large numbers limit, as $\Omega\to\infty$, of the Markov process governed by the master equation [35, 34]. The Linear Noise Approximation (LNA), derived using the system size expansion method, is a stochastic model that provides even more rapid simulation than the aforementioned stochastic methods as well as being the only model in this field that allows analytical expressions of probability distributions of populations at arbitrary time-points. This makes the LNA a powerful tool for sensitivity analysis and statistical inference [46, 47, 61, 48, 33, 12, 22, 55]. However, the important drawback of using the LNA is its inability to approximate the long-term behavior of non-linear population dynamics. Indeed, when the system dynamics are non-linear [69], predictions of the state of the system at time $t$ given the state at an earlier time $s$, $0<s<t$, become inaccurate as tst-s becomes large.
@@
For any given system, we can qualitatively examine the dynamics of its paths and identify different phases in the populations’ evolution—for instance their peaks and troughs. If we were to examine the distribution of entire stochastic paths generated by the SSA in a finite number of time-points then we find that, the phases of such paths become out of sync with what the LNA predicts their phases to be. Hence the LNA makes incorrect predictions of the dynamics thereafter. In [46], it was shown that in the specific setting where the macroscopic ($\Omega\to\infty$) deterministic approximation has an attractive limit cycle solution, correcting these drifts in phase predicted by the LNA is sufficient to produce simulations that are nearly identical to those produced by the SSA. This approach enables the construction of both fast and accurate algorithms for stochastic simulation, sensitivity analysis, and statistical inference of this class of non-linear systems [46].
The main results of this paper are the following. Firstly, we develop a novel framework for constructing LNA-based models for non-linear dynamical systems. We adopt the term phase corrected LNA (pcLNA)—originally introduced in [46]—to describe this stochastic modelling approach, which evolves the system using LNA equations while periodically adjusting the phase to mitigate phase drift. Fundamental to the pcLNA approach is a mapping G{G} that computes the phase of any stochastic state allowing for any phase drift to be corrected. We show that, for any parameter-dependent system whose macroscopic reaction-rate equation admits a non-hyperbolic equilibrium for some parameter value, tools from dynamical systems theory can be used to explicitly construct the phase-correcting map G{G}. Notably, by combining results from [51, 62] with Itô’s Lemma [38], we derive a novel decomposition of the LNA stochastic process into coordinates exhibiting non-linear, persistent dynamics and coordinates with transient dynamics. This decomposition suggests that a common phase definition can be used for classes of stochastic systems with qualitatively similar behavior. Using these findings, we provide a step-by-step procedure for defining a phase-correcting map and a general algorithm to be used for stochastic simulation.
Secondly, we use this framework to provide readily applicable algorithms that can accurately simulate long-term stochastic dynamics for arguably the two most abundant non-linear phenomena: oscillations and bi-stability. We demonstrate this on several systems including the oscillatory systems in [39, 71, 4], and bi-stable systems in [14, 2, 21]. Here the pcLNA model offers accurate approximations of the master equation for long-term simulation whilst substantially reducing the simulation time by several orders of magnitude. We also propose a fast method to predict the speed of SSA simulations, and therefore assess whether a faster simulation method is necessary.
We emphasize that with further development on defining explicit phase-correcting maps $G$, the pcLNA framework introduced here could be used to simulate other class of non-linear dynamical systems with computational efficiency and accuracy. We also note that a key implication of the speed and accuracy of pcLNA is that it enables the approach to be used for computationally intensive tasks such as sensitivity analysis, inference, and prediction for large, non-linear systems.
@@
While the approach developed in [46] demonstrated that phase correction significantly improves the long-term accuracy of the LNA in oscillatory systems with stable limit cycles, it was limited to this specific dynamics. In contrast, the framework introduced in this paper generalises the concept of phase correction by introducing centre manifold theory to define and correct phase drifts to a broad class of systems only assumed to have a non-hyperbolic equilibrium, which includes the abundant classes of oscillatory and bi-stable systems.
@@
We begin in Section 2 with a review of stochastic population dynamics evolving through reaction networks (Section 2.1) and the system size expansion (Section 2.2), and then introduce the Linear Noise Approximation (LNA) in Section 2.3. We conclude the section by discussing the LNA’s loss of accuracy in non-linear systems. In Section 3, we introduce the phase corrected Linear Noise Approximation framework. To understand the LNA’s failure we define the notion of phase in the context of stochastic reaction networks (Section 3.1) before presenting the general phase corrected LNA (pcLNA) simulation algorithm in Section 3.2. We follow this by a discussion on how to define the phase-correcting map in Section 3.3. In Section 4, we apply our framework to reaction networks whose macroscopic reaction rate equation undergoes a Hopf bifurcation and develop the pcLNA algorithm specific to these systems. Section 5 focuses on bi-stable systems and constructs the corresponding pcLNA algorithm. We validate our framework through numerical investigations in Section 6, and conclude with a discussion of our findings and future directions in Section 7. Additional numerical investigations demonstrating the accuracy of the pcLNA modelling techniques on other reaction networks, as well as supporting technical results and derivations, are provided in the Supporting Information (SI).
@@
2 Population dynamics
We begin by reviewing, in Sections 2.1-2.3, the mathematical formulation of population dynamics evolving through reaction networks. Our discussion is restricted to the concepts needed for the development of our methodology; for a more comprehensive review, see [1, 45, 72]. We describe this formulation within the context of molecular biology, although analogous terminology can be applied in other fields such as epidemiology and ecology (see, for instance, [22, 72] and Figure 2).
2.1 Reaction networks
A system of $n$ different molecular species, $M_{1},M_{2},\dots,M_{n}$ has state vector, ${\bm{Y}}(t)=(Y_{1}(t),\ldots,Y_{n}(t))^{\top}$ where $Y_{i}(t)$, $i\in\{1,\ldots,n\}$, denotes the number of molecules (i.e. the population) of species $M_{i}$ at time $t$ and $\top$ denotes the transpose. These molecules undergo reactions $R_{j}$, $j\in\{1,\dots,r\}$, such as transcription, translation, degradation and translocation, which change the number of molecules of each species,
@@ 2.1 Reaction networks
Here, the non-negative integer coefficients $k_{ij}$ and $k_{ij}^{\prime}$ denote the number of molecules of species $M_{i}$ involved as reactants and products, respectively, in reaction $R_{j}$. If $R_{j}$ occurs at time $t\geq 0$, ${\bm{Y}}(t)$ jumps to a new state ${\bm{Y}}(t)+{\bm{\nu}}_{j}$, where the integer transition vectors ${\bm{\nu}}_{j}=(k_{1j}^{\prime}-k_{1j},\dots,k_{nj}^{\prime}-k_{nj})^{\top}$ are often called stoichiometric. Under the assumption that these reactions occur in a well-mixed solution [17], we can derive the propensity functions $\pi_{j}:\mathbb{R}^{n}\to[0,\infty)$ for each $j\in\{1,\dots,r\}$ with the probability of one $R_{j}$ reaction occurring in the next infinitesimal time interval $[t,t+dt)$
@@
The propensity functions $\pi_{j}$ along with the stoichiometric vector ${\bm{\nu}}_{j}$, $j=1,\dots,r$, completely specify the evolution of the reaction network.
[…]
These conditions give rise to a time-inhomogeneous, discrete state-space Markov process $\{{\bm{Y}}(t):t\geq 0\}$. The Kolmogorov forward equation [18] describing the evolution of the probability distribution, $P({\bm{y}},t):=P({\bm{Y}}(t)={\bm{y}}\mid{\bm{Y}}(0)={\bm{y}}_{0})$, of the stochastic process $\{{\bm{Y}}(t):t\geq 0\}$,
$\frac{\partial P({\bm{y}},t)}{\partial t}=\sum_{j=1}^{r}\pi_{j}({\bm{y}}-{\bm{\nu}}_{j})P({\bm{y}}-{\bm{\nu}}_{j},t)-\sum_{j=1}^{r}\pi_{j}({\bm{y}})P({\bm{y}},t),$ (3)
is often referred to as the (chemical) master equation. Despite the equation only being analytically tractable for few cases, exact trajectories can be generated using Gillespie’s stochastic simulation algorithm (SSA) [17]. We provide details of this in SI (Section S1).
Equivalently, the process can be represented via the random time change representation (RTC) [1]
where $N_{j}$, $j\in\{1,\dots,r\}$, are independent, unit-rate Poisson processes corresponding to reaction $R_{j}$. For a given trajectory ${\bm{y}}(s)$, $s\in[0,t)$, the random variables $N_{j}\left(\int_{0}^{t}\pi_{j}({\bm{y}}(s))ds\right)$ are independent and have Poisson distribution with mean $\int_{0}^{t}\pi_{j}({\bm{y}}(s))ds$.
Example (SIR model)
The SIR model used to describe population dynamics in infectious disease outbreaks is an example of a reaction network. We formulate it here to provide a concrete instance of the general framework introduced above (see also Figure 2(A-B) for a summary). In the SIR network, there are three populations, say $M_{1},M_{2},M_{3}$, corresponding to the Susceptible (S), Infected (I), and Recovered (R) individuals. Two reactions can occur—infection ($R_{c}$) and recovery ($R_{2}$)—which can be represented as
@@
where $c_{1}$ and $c_{2}$ are the reaction rate constants for the two reactions. The state vector ${\bm{Y}}=(Y_{1},Y_{2},Y_{3})^{\top}$ represents the three populations $M_{1}$ (S), $M_{2}$ (I), and $M_{3}$ (R). The propensity functions are
@@
with corresponding stoichiometry vectors ${\bm{v}}_{1}=(-1,+1,0)^{\top}$ and ${\bm{v}}_{2}=(0,-1,+1)^{\top}$. This can be used to formulate Equations (3) and (4) for this system.